When Rs 165 is divided in the ratio 2 : 4 : 5 , what is the difference between the largest and the smallest share?

Answer by rothauserc(4718)

2x +4x +5x = 165
11x = 165
x = 15
the three shares are 30, 60, 75
note that 30 + 60 + 75 = 165 and
75 - 30 = 45
therefore,
The difference between the largest and the smallest share is Rs 45